Assume that an electron is located somewhere within a region of atomic size. Estimate the minimum uncertainty in its momentum by assuming that this uncertainty is comparable with its average momentum. Estimate the average kinetic energy of the electron.
1
Expert's answer
2012-08-28T10:33:03-0400
sigmax*sigmap >=h/2

sigma
x=10^-8

sigma p=p=3*10^-19

E= p^2/ (2m) = 9*10^(-38)/(2*9*10^-34) = 0.5*10^-4